# Statistical Inference for Policy Evaluation with Temporal Difference Learning

### v3

@@ Abstract
We investigate the statistical properties of Temporal Difference (TD) learning with Polyak-Ruppert averaging, arguably one of the most widely used algorithms in reinforcement learning, for the task of estimating the parameters of the optimal linear approximation to the value function. We make three significant contributions that improve the current state-of-the-art results: (i) we derive sharper high probability convergence guarantee that depend explicitly on the asymptotic variance and hold under weaker conditions than those normally assumed; (ii) we establish refined high-dimensional Berry-Esseen bounds over the class of convex sets, achieving faster rates than those previously established in the literature, and (iii) we propose and analyze a novel, computationally efficient online plug-in estimator of the asymptotic covariance matrix. These results enable the construction of confidence regions and simultaneous confidence intervals for the linear parameters of the value function approximation, with guaranteed finite-sample coverage. We demonstrate the applicability of our theoretical findings through numerical experiments.
@@ 1 Introduction
A critical task in RL is to evaluate with confidence the quality of a given policy, measured by its value function. As RL models are based on Markov Decision Processes (MDPs), statistical inference must accommodate the online nature of the sampling mechanism. TD learning (see, e.g., Sutton, 1988), an iterative procedure that can be framed as an instance of stochastic approximation (SA), is arguably the most widely used and effective algorithm for policy evaluation in RL. The last several years have witnessed a flurry of activities within the statistical machine learning community to derive finite sample consistency rates of TD iterates, leading to deep results and a much improved understanding of the performance and limitations TD learning. See, e.g., Khamaru et al. (2020); Li et al. (2024); Bhandari et al. (2021); Dalal et al. (2018); Srikant and Ying (2019); Lakshminarayanan and Szepesvari (2018); Mou et al. (2020); Durmus et al. (2022); Li et al. (2024); Samsonov et al. (2024); Prashanth et al. (2013); Borkar et al. (2025) for a partial and growing list of relevant references. Though for the tabular case (finite state and action spaces) the TD algorithm is known to be minimax rate optimal Li et al. (2024), its properties – and, in particular, its sample complexity – in more complex and high-dimensional settings have not been fully understood.
@@
Li et al. (2024) $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log^{\frac{1}{2}}\frac{T}{\delta}\right)$ $\eta\sqrt{\log\frac{T}{\delta}}\leq\tilde{C}$
Samsonov et al. (2024) $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log\frac{T}{\delta}\right)$ $\eta\log\frac{T}{\delta}\leq\tilde{C}$
Theorem 3.1 $\eta_{t}=\eta_{0}t^{-\alpha},\alpha\in\left(\frac{1}{2},1\right)$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\sqrt{\log\frac{d}{\delta}}\right)$ $\eta_{0}\sqrt{\log\frac{1}{\delta\eta_{0}}}\leq\tilde{C}$
Table 1: Comparisons between our results on TD convergence rate with prior work. The quantity $\tilde{C}$, whose value may change at each occurrence, may depend on problem-related parameters other than $T$ and $\delta$.
In a more specialized line of work, concerned with establishing distributional approximations of the TD iterates, it has been shown that, using Polyak-Ruppert averaging (Polyak and Juditsky, 1992; Ruppert, 1998), the TD estimation error converges to a Gaussian distribution in the large sample limit; see, e.g., Fort (2015); Mou et al. (2020); Li et al. (2023). While these results provide a formal justification for the validity of statistical inference, the guarantees provided in those contributions are non-quantitative, in the sense that they are of asymptotic nature and hold only in fixed dimensions. As a result, they do not adequately capture the dependence on the sample size and other problem-related quantities, most notably the dimensionality, all of which affect the accuracy of statistical inference. In this paper, we partially address this gap. We focus on TD learning with Polyak-Ruppert averaging for estimating the coefficients of the best linear approximation of the value function, assuming i.i.d. samples. We make the following theoretical and methodological contributions.
@@
    Sharper Berry-Esseen bounds. We derive a new high-dimensional Berry-Esseen bound for the estimation error over the class of convex sets and using an estimated variance, achieving an $O(T^{-\frac{1}{3}})$ rate as a function of the number of iterations, which appears to be the fastest in the literature. While completing this work, we became aware of the recent related paper by Samsonov et al. (2024), which, under the same settings, derived a high-dimensional Berry-Esseen bound for linear SA with Polyak-Ruppert averaging, obtaining a slower rate, of order $O(T^{-\frac{1}{4}})$. See Section 3.2 for detailed comparisons.
@@
    Novel, computationally efficient variance estimator. We develop a computationally efficient online algorithm to construct an estimator $\widehat{\bm{\Lambda}}_{T}$ of the asymptotic variance $\bm{\Lambda}^{\star}$. We also show that the total variation distance between the (computable) distribution $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ and the asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ vanishes with high probability at a rate $O(T^{-\frac{1}{3}})$.
@@
We verify our theoretical findings with some numerical simulations, discussed in the appendix. As an illustration, Figure 1 shows how the empirical distribution of TD estimation error in $1,000$ independent trials (as shown in blue histograms), scaled by the square-root of the number of iterations $T$, converges to its asymptotic Gaussian distribution (shown by the red curves).
The i.i.d. condition. Our results are derived under the simplifying but standard assumption of i.i.d. data, which has been widely used in the recent literature on statistical analyses of TD learning, SA and stochastic gradient descent (see, e.g. Sheshukova et al., 2025). Analogous results, obtained using different techniques and yielding slower rates, have been recently obtained for the more challenging case of Markovian data: see Srikant (2024) and Wu et al. (2025).
Notation. We use boldface small letters to denote vectors and capital letters to denote matrices. For a vector $\bm{x}$, $\|\bm{x}\|_{2}$ incates its Euclidean norm; for a matrix $\bm{M}$, we use $\mathsf{Tr}(\bm{M})$, $\mathsf{det}(\bm{M})$, $\|\bm{M}\|_{\mathsf{F}}$ and $\|\bm{M}\|$ to denote its trace, determinant, Frobenius and spectral norm, respectively. For a symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we use $\lambda_{\max}(\bm{M})=\lambda_{1}(\bm{M})\geq\ldots\lambda_{d}(\bm{M})=\lambda_{\min}(\bm{M})$ to denote its eigenvalues and $\mathrm{cond}(\bm{M})$ to indicate its condition number. For a vector $\bm{x}\in\mathbb{R}^{d}$ and a symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we let $\|\bm{x}\|_{\bm{M}}=\sqrt{\bm{x}^{\top}\bm{Mx}}$. We will use the notation $\tilde{C}$ to express a quantity independent of $T$, the number of iterations/sample size, but possibly dependent on other problem-related parameters. Similarly, for positive sequences $\{f_{t}\}_{t\in\mathbb{N}}$ and $\{g_{t}\}_{t\in\mathbb{N}}$, we write $f_{t}\lesssim g_{t}$, or $f_{t}=O(g_{t})$, to signify that $f_{t}\leq\tilde{C}g_{t}$, for some $C>0$ independent of $t$. The convex distance between two probability distributions on $\mathbb{R}^{d}$, say $P$ and $Q$, is defined as
@@ 2 Background
An infinite-horizon Markov Decision Process (MDP) (Bertsekas, 2017; Sutton and Barto, 2018) is the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ and $\mathcal{A}$ represent the state space of the environment and the action space available to an agent, respectively, both assumed for simplicity to be finite (though our results hold for countable spaces). The function $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is a transition kernel that associates to any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ the probability distribution $P(\cdot\mid s,a)$ of the next state of the environment when the agent executes action $a$ while the environment is in state $s$. The reward function $r:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ measures the immediate reward to the agent after taking action $a$ at state $s$; finally, $\gamma\in[0,1)$ is the discount factor, used to assess the impact of future rewards.
In choosing their actions, the agent follows a randomized action selection rule, or policy, represented as a (time-invariant) conditional probability distribution $\pi:\mathcal{S}\mapsto\Delta(\mathcal{A})$ over the action space given the state of the environment. For any policy $\pi$, its associated reward function $r:\mathcal{S}\mapsto[0,1]$ is the expected reward of the next action selected according to the policy $\pi$, i.e. $r(s):=\mathbb{E}_{a\sim\pi(\cdot\mid s)}[r(s,a)],$ for $s\in\mathcal{S}.$ Any policy $\pi$ induces a transition kernel $P$, representing the probability distribution of the next state given the current state $s$ and the policy $\pi$, given by $P(s^{\prime}\mid s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}P(s^{\prime}\mid s,a)$ for $s,s^{\prime}\in\mathcal{S}.$ We assume throughout that there exists a unique stationary distribution $\mu$ for the transition matrix ${P}$; that is, $\mu(i)=\mathbb{P}_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}(s^{\prime}=i)$, $i\in\mathcal{S}$.
@@
$V(s)=r(s)+\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}V(s^{\prime})=:\mathcal{T}V(s).$ (1)
@@
In most practical scenarios, the value function is very complex and impossible to compute due to the prohibitively large size of the state space $\mathcal{S}$ A common way to overcome this issue is to consider a parametric approximation of $V$ through a linear function, which is both computationally tractable and amenable to statistical estimation. In detail, for a given feature map $\bm{\phi}:\mathcal{S}\mapsto\mathbb{R}^{d}$, mapping each state in $\mathcal{S}$ to a centered feature vector in $\mathbb{R}^{d}$, where $d$ is typically small(er than $|\mathcal{S}|$), a linear value function approximation is given by
@@
where ${\bm{\theta}}\in\mathbb{R}^{d}$ is a vector of linear coefficients. Throughout this paper, we assume that the Eculidean norm of the features are uniformly bounded, i.e., $\|\bm{\phi}(s)\|_{2}\leq 1$ for all $s\in\mathcal{S}$.
The best approximating linear function is the one parametrized by the coefficient $\bm{\theta}^{\star}$ solving the projected Bellman equation (Tsitsiklis and Van Roy, 1997) $V_{\bm{\theta}^{\star}}=\Pi\mathcal{T}V_{\bm{\theta}^{\star}},$ with $\Pi$ the orthogonal projection onto the subspace of functions that can be represented by linear combinations of the features. See, e.g., Duan et al. (2021). In turn, this (unique!) solution is given by
@@
$\bm{A}:=\mathop{\mathbb{E}}\limits_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}\left[\bm{\phi}(s)\left(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})\right)^{\top}\right]\in\mathbb{R}^{d\times d}\quad\text{and}\quad\bm{b}:=\mathop{\mathbb{E}}\limits_{s\sim\mu}\left[\bm{\phi}(s)r(s)\right]\in\mathbb{R}^{d}.$ (4)
@@
$\bm{\Sigma}:=\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}^{\top}(s)],\quad\lambda_{0}:=\lambda_{\min}(\bm{\Sigma}),\quad\text{and}\quad\lambda_{\Sigma}:=\lambda_{\max}(\bm{\Sigma})$ (5)
be the feature covariance matrix and its smallest and largest eigenvalues, respectively.
@@
While in most real-world RL scenarios the state-action-reward tuples follow a Markov decision process, many of the most recent theoretical contributions – including those delivering the sharpest guarantees listed in Table 1 and Samsonov et al. (2024) – have made the simplifying but technically very convenient assumption of $i.i.d.$ data. Here we follow the same convention and analyze the properties of the TD learning algorithm applied to a sequence of $T$ $i.i.d.$ state-action-reward tuples of the form
@@
where $(s_{1},\ldots,s_{T})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu$, the stationary distribution, and, for eact $t$, $a_{t}\sim\pi(s_{t}),r_{t}=r(s_{t},a_{t})$ and $s_{t}^{\prime}\sim P(\cdot\mid s_{t},a_{t})$.
@@
This estimator is an instance of the Polyak-Ruppert averaging scheme (Polyak and Juditsky, 1992; Ruppert, 1998), a stochastic approximation methodology that provably reduces the variability and accelerates convergence. Furthermore, with a proper choice of the stepsizes, the averaged iterates converge to a Gaussian distribution, thus enabling statistical inference. To date, the best Berry-Esseen bounds rates for stochastic approximation methods and TD learning with Polyak-Ruppert averaging are due to Samsonov et al. (2024), which in turn are based on general non-linear Berry-Esseeen bounds by Shao and Zhang (2022). In this paper we improve on these results by deriving novel and sharper high-probability and Berry-Esseen bounds for the estimation error of the TD algorithm

### v4

@@ Abstract
We investigate the statistical properties of Temporal Difference (TD) learning with Polyak-Ruppert averaging, arguably one of the most widely used algorithms in reinforcement learning, for the task of estimating the parameters of the optimal linear approximation to the value function. Assuming independent samples, we make three theoretical contributions that improve upon the current state-of-the-art results: (i) we derive sharper high probability convergence guarantees that depend explicitly on the asymptotic variance and hold under weaker conditions than those adopted in the literature; (ii) we establish refined high-dimensional Berry-Esseen bounds over the class of convex sets, achieving faster rates than the best known results, and (iii) we propose and analyze a novel, computationally efficient online plug-in estimator of the asymptotic covariance matrix. These results enable the construction of confidence regions and simultaneous confidence intervals for the linear parameters of the value function approximation, with guaranteed finite-sample coverage. We demonstrate the applicability of our theoretical findings through numerical experiments.
@@ 1 INTRODUCTION
A critical task in RL is to evaluate with confidence the quality of a given policy, measured by its value function. As RL models are based on Markov Decision Processes (MDPs), statistical inference must accommodate the online nature of the sampling mechanism. TD learning (see, e.g., Sutton,, 1988), an iterative procedure that can be framed as an instance of stochastic approximation (SA), is arguably the most widely used and effective algorithm for policy evaluation in RL. The last several years have witnessed a flurry of activities within the statistical machine learning community to derive finite sample consistency rates of TD iterates, leading to deep results and a much improved understanding of the performance and limitations TD learning. See, e.g., Khamaru et al., (2020); Li et al., 2024a ; Bhandari et al., (2021); Dalal et al., (2018); Srikant and Ying, (2019); Lakshminarayanan and Szepesvari, (2018); Mou et al., (2020); Durmus et al., (2022); Li et al., 2024b ; Samsonov et al., 2024b ; Prashanth et al., (2013); Borkar et al., (2025) for a partial and growing list of relevant references. Though for the tabular case (finite state and action spaces) the TD algorithm is known to be minimax rate optimal Li et al., 2024a , its properties – and, in particular, its sample complexity – in more complex and high-dimensional settings have not been fully understood.
@@
Li et al., 2024b $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log^{\frac{1}{2}}\frac{T}{\delta}\right)$ $\eta\sqrt{\log\frac{T}{\delta}}\leq\widetilde{C}$
Samsonov et al., 2024b $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log\frac{T}{\delta}\right)$ $\eta\log\frac{T}{\delta}\leq\widetilde{C}$
Theorem 3.1 $\eta_{t}=\eta_{0}t^{-\alpha},\alpha\in\left(\frac{1}{2},1\right)$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\sqrt{\log\frac{1}{\delta}}\right)$ $\eta_{0}\sqrt{\log\frac{1}{\delta\eta_{0}}}\leq\widetilde{C}$
Table 1: Comparisons between our results on TD convergence rate with prior work. The quantity $\widetilde{C}$, whose value may change at each occurrence, may depend on problem-related parameters other than $T$ and $\delta$. The conditions required in Theorem 3.1 are weaker than those in Samsonov et al., 2024b and Li et al., 2024b while delivering the same rates; see discussion in the paragraph below Theorem 3.1.
In a more specialized line of work, concerned with establishing distributional approximations of the TD iterates, it has been shown that, using Polyak-Ruppert averaging (Polyak and Juditsky,, 1992; Ruppert,, 1998), the TD estimation error converges to a Gaussian distribution in the large sample limit; see, e.g., Fort, (2015); Mou et al., (2020); Li et al., (2023). While these results provide a formal justification for the validity of statistical inference, the guarantees from those contributions are non-quantitative, in the sense that they are of asymptotic nature and hold only in fixed dimensions. As a result, they do not adequately capture the dependence on the sample size and other problem-related quantities, most notably the dimensionality, all of which affect the accuracy of statistical inference. In this paper, we partially address this gap. We focus on TD learning with Polyak-Ruppert averaging for estimating the coefficients of the best linear approximation of the value function, assuming i.i.d. samples. We make the following theoretical and methodological contributions.
@@
    Sharper Berry-Esseen bounds. We derive a new high-dimensional Berry-Esseen bound for the estimation error over the class of convex sets and using an estimated variance, achieving an $O(T^{-\frac{1}{3}})$ rate as a function of the number of iterations, which appears to be the fastest in the literature. While completing this work, we became aware of the recent related paper by Samsonov et al., 2024a , which, under the same settings, derived a high-dimensional Berry-Esseen bound for linear SA with Polyak-Ruppert averaging, obtaining the slower rate $O(T^{-\frac{1}{4}})$. See Section 3.2 for a discussion and detailed comparisons.
@@
    Novel, computationally efficient variance estimator. We develop a computationally efficient online algorithm to construct an estimator $\widehat{\bm{\Lambda}}_{T}$ of the asymptotic variance $\bm{\Lambda}^{\star}$. We also show that the total variation distance between the (computable) distribution $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ and the asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ vanishes with high probability at a rate $O(T^{-\frac{1}{3}})$, matching the order of our Berry-Esseen rate. As a result, estimating the variance, a crucial step for constructing confidence sets, does not led to worse rates.
@@
We verify our theoretical findings with numerical simulations, discussed in the appendix. As an illustration, Figure 1 shows how the empirical distribution of TD estimation error in $1,000$ independent trials (as shown in blue histograms), scaled by the square-root of the number of iterations $T$, converges to its asymptotic Gaussian distribution (shown by the red curves).
The i.i.d. condition. Our results are derived under the standard assumption of i.i.d. data, which has been widely used in the recent literature on statistical analyses of TD learning, SA and stochastic gradient descent (see, e.g. Sheshukova et al.,, 2025). While unrealistic, the theoretical investigation of this simplified scenario has proved challenging; the present work addresses some of the gap still existing in the literature and provide state-of-the art rates. We mentioon that analogous results, obtained using different techniques and yielding slower rates, have been recently obtained for the more challenging case of Markovian data: see Srikant, (2024) and Wu et al., (2025).
Notation. We use boldface small letters to denote vectors and capital letters to denote matrices. For a vector $\bm{x}$, $\|\bm{x}\|_{2}$ incates its Euclidean norm; for a matrix $\bm{M}$, we use $\mathsf{Tr}(\bm{M})$, $\mathsf{det}(\bm{M})$, $\|\bm{M}\|_{\mathsf{F}}$ and $\|\bm{M}\|$ to denote its trace, determinant, Frobenius and spectral norm, respectively. For a symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we use $\lambda_{\max}(\bm{M})=\lambda_{1}(\bm{M})\geq\ldots\lambda_{d}(\bm{M})=\lambda_{\min}(\bm{M})$ to denote its eigenvalues and $\mathrm{cond}(\bm{M})$ to indicate its condition number. For a vector $\bm{x}\in\mathbb{R}^{d}$ and a symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we let $\|\bm{x}\|_{\bm{M}}=\sqrt{\bm{x}^{\top}\bm{Mx}}$. We will use the notation $\widetilde{C}$ to express a quantity independent of $T$, the number of iterations/sample size, but possibly dependent on other problem-related parameters. Similarly, for positive sequences $\{f_{t}\}_{t\in\mathbb{N}}$ and $\{g_{t}\}_{t\in\mathbb{N}}$, we write $f_{t}\lesssim g_{t}$, or $f_{t}=O(g_{t})$, to signify that $f_{t}\leq\widetilde{C}g_{t}$, for some $C>0$ independent of $t$. The convex distance between two probability distributions on $\mathbb{R}^{d}$, say $P$ and $Q$, is defined as
@@ 2 BACKGROUND
An infinite-horizon discounted Markov Decision Process (MDP) (Bertsekas,, 2017; Sutton and Barto,, 2018) is the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ and $\mathcal{A}$ represent the state space of the environment and the action space available to an agent, respectively, both assumed for simplicity to be finite (though our results hold for countable spaces). The function $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is a transition kernel that associates to any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ the probability distribution $P(\cdot\mid s,a)$ of the next state of the environment when the agent executes action $a$ while the environment is in state $s$. The reward function $r:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ measures the immediate reward to the agent after taking action $a$ at state $s$; finally, $\gamma\in[0,1)$ is the discount factor, used to assess the impact of future rewards.
In choosing their actions, the agent follows a randomized action selection rule, or policy, represented as a (time-invariant) conditional probability distribution $\pi:\mathcal{S}\mapsto\Delta(\mathcal{A})$ over the action space given the state of the environment. For any policy $\pi$, its associated reward function $r:\mathcal{S}\mapsto[0,1]$ is the expected reward of the next action selected according to the policy $\pi$, i.e. $r(s):=\mathbb{E}_{a\sim\pi(\cdot\mid s)}[r(s,a)],$ for $s\in\mathcal{S}.$ Any policy $\pi$ induces a transition kernel $P$, representing the probability distribution of the next state given the current state $s$ and the policy $\pi$, given by $P(s^{\prime}\mid s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}P(s^{\prime}\mid s,a)$ for $s,s^{\prime}\in\mathcal{S}.$ We assume throughout that there exists a unique stationary distribution $\mu$ for the transition matrix P{P}; that is, $\mu(i)=\mathbb{P}_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}(s^{\prime}=i)$, $i\in\mathcal{S}$.
@@
$V(s)=r(s)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}V(s^{\prime})=:\mathcal{T}V(s).$ (1)
@@
In most practical scenarios, the value function is very complex and impossible to compute due to the prohibitively large size of the state space $\mathcal{S}$. A common way to overcome this issue is to consider a parametric approximation of $V$ through a linear function, which is both computationally tractable and amenable to statistical estimation. In detail, for a given feature map $\bm{\phi}:\mathcal{S}\mapsto\mathbb{R}^{d}$, mapping each state in $\mathcal{S}$ to a centered feature vector in $\mathbb{R}^{d}$, where $d$ is typically smaller than $|\mathcal{S}|$, a linear value function approximation is given by
@@
where ${\bm{\theta}}\in\mathbb{R}^{d}$ is a vector of linear coefficients. Throughout this paper, we assume that the Euclidean norm of the features are uniformly bounded, i.e., $\|\bm{\phi}(s)\|_{2}\leq 1$ for all $s\in\mathcal{S}$.
The best approximating linear function is the one parametrized by the coefficient $\bm{\theta}^{\star}$ solving the projected Bellman equation (Tsitsiklis and Van Roy,, 1997) $V_{\bm{\theta}^{\star}}=\Pi\mathcal{T}V_{\bm{\theta}^{\star}},$ with $\Pi$ the orthogonal projection onto the subspace of functions that can be represented by linear combinations of the features. See, e.g., Duan et al., (2021). In turn, this (unique!) solution is given by
@@
$\bm{A}:=\mathop{\mathbb{E}}\limits_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}\left[\bm{\phi}(s)\left(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})\right)^{\top}\right]\in\mathbb{R}^{d\times d}$ (4)
$\text{and}\quad\bm{b}:=\mathop{\mathbb{E}}\limits_{s\sim\mu}\left[\bm{\phi}(s)r(s)\right]\in\mathbb{R}^{d}.$ (5)
@@
$\bm{\Sigma}:=\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}^{\top}(s)]$ (6)
be the feature covariance matrix and use $\lambda_{0}$, $\lambda_{\Sigma}$ to denote its smallest and largest eigenvalues, respectively.
@@
While in most real-world RL scenarios the state-action-reward tuples follow a Markov decision process, many of the most recent theoretical contributions – including those delivering the sharpest guarantees listed in Table 1 and Samsonov et al., 2024a – have made the simplifying but technically very convenient assumption of $i.i.d.$ data. Here we follow the same convention and analyze the properties of the TD learning algorithm applied to a sequence of $T$ $i.i.d.$ state-action-reward tuples of the form
@@
where $(s_{1},\ldots,s_{T})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu$, the stationary distribution, and, for each $t$, $a_{t}\sim\pi(s_{t}),r_{t}=r(s_{t},a_{t})$ and $s_{t}^{\prime}\sim P(\cdot\mid s_{t},a_{t})$.
@@
This estimator is an instance of the Polyak-Ruppert averaging scheme (Polyak and Juditsky,, 1992; Ruppert,, 1998), a stochastic approximation methodology that provably reduces the variability and accelerates convergence. Furthermore, with a proper choice of the stepsizes, the averaged iterates converge to a Gaussian distribution, thus enabling statistical inference. To date, the best Berry-Esseen bounds rates for stochastic approximation methods and TD learning with Polyak-Ruppert averaging are due to Samsonov et al., 2024a , which in turn are based on general non-linear Berry-Esseen bounds by Shao and Zhang, (2022). In this paper we improve on these results by deriving novel and sharper high-probability and Berry-Esseen bounds for the estimation error of the TD algorithm